OpenStudy (anonymous):

These are the heights of Natasha and her four friends: 5 feet, 2 inches; 5 feet, 3 inches; 5 feet, 7 inches; 5 feet, 9 inches; 6 feet, 0 inches In which situation would the mean of the heights go up the most? Select one of the options below as your answer: A. You add 2 friends that are 5 feet, 8 inches. B. You add 1 friend that is 6 feet, 5 inches and 1 friend that is 5 feet, 3 inches. C. You add 1 friend that is 5 feet, 10 inches. D. You add 4 friends that are 5 feet, 9 inches.

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

The "mean" means the average. I'm not seeing an easy, straight-forward way to find this answer but somebody else might. The way I'd do it is just trial and error. First convert everyone's heights to inches so it's all in the same units so it's easier to calculate. There are 12 inches in 1 foot so: *1st person: 5ft,2in... 5*12=60in plus those 2in means a height of 62in. *2nd person: 5ft,3in... 5*12=60in plus those 3in means a height of 63in. *3rd person: 5ft, 7in... 5*12=60in plus those 7in means a height of 67in. ...and so on for the other two. Then add all those heights up so you get SOME TOTAL NUMBER. Take that number and divide it by the total number of people (5) to get the average. Then go through each of your choices and see how many inches they gave you. For example, (A) gives you 2 friends who are 5*12=60 plus 8in ... 68 inches. 2 * 68 = 136in. Add that 136in to the SOME TOTAL NUMBER from the beginning and divide by the total number of people.. the total number of people now went up from 5 to 5+2 which is 7, so 126in plus SOME TOTAL NUMBER all divided by 7 gives the average. Do the same for B, C, D and see which one gives the highest average.

OpenStudy (anonymous):

For B, the total number of people would be those original 5 plus two friends again, so 7 again. For C, the total number of people would be the original 5 plus one friend you add so 5+1 = 6. For D, the total number of people would be 5 + 4 = 9. That should be enough setup to get you started.

OpenStudy (shane_b):

I agree that's probably the best way to do it...although you can almost guess that it's going to be D based on the other answers.